OpenStudy (anonymous):

lets see who is more intelligent will get a medal from me 3+3x3-3+3=?

OpenStudy (briannabug):

3x3=9 9-3=6+3+3=12

OpenStudy (anonymous):

which rule have u used because i used DMAS rule an im getting 6

OpenStudy (unklerhaukus):

\[3+3\times3-3+3=((3+(3\times3))-3)+3\]

OpenStudy (unklerhaukus):

\[3+3\times3-3+3\\=((3+(3\times3))-3)+3\\ =((3+9)-3)+3\\ =(12-3)+3\\ =9+3\\ = \]

OpenStudy (briannabug):

THANKS FOR THE MEADAL